OpenStudy (anonymous):

simplify: sin^3 (t) + cos^3 (t) _______________________ sin (t) + cos (t)

OpenStudy (anonymous):

the dotted line means dividing

OpenStudy (anonymous):

are you there?

OpenStudy (rizags):

yea im working on it

OpenStudy (anonymous):

k thanks

OpenStudy (rizags):

do you want just answer or explanation too?

OpenStudy (anonymous):

explanation plz

OpenStudy (anonymous):

hello?

OpenStudy (rizags):

\[(a ^{3}+b ^{3})=(a+b)(a ^{2}-ab+b^{2})\]

OpenStudy (anonymous):

oh, a difference of cubes...

OpenStudy (rizags):

sooOOOOO.... \[\sin ^{3}t+\cos ^{3}t=(\sin t +\cos t)(\sin ^{2}t-\sin t \cos t+\cos ^{2}t)\]

OpenStudy (rizags):

divide thru the sin t + cos t

OpenStudy (rizags):

then you get \[\sin ^{2}t+\cos ^{2}t - \sin t \cos t= 1- \sin t \cos t\]

OpenStudy (rizags):

thats the final answer i think

OpenStudy (anonymous):

yes, then we know that sin squared plus cos squared equals one

OpenStudy (anonymous):

yep i see now, thanks!

OpenStudy (rizags):

medal plz

OpenStudy (rizags):

ty

OpenStudy (anonymous):

deserved!